Suppose Compco Systems pays no dividends but spent $4.92 billion on share repurchases last year. If​ Compco's equity cost of capital is 12.8%​, and if the amount spent on repurchases is expected to grow by 8.4% per​ year, estimate​ Compco's market capitalization. If Compco has 6.6 billion shares​ outstanding, to what stock price does this​ correspond?​Compco's market capitalization will be? billion. (Round to two decimal​ places.)

Please show steps to reach final answer.
Answer #1

a). Total Payout = Payout Last Year * (1 + g) = $4.92 billion * (1 + 0.084) = $5.33 billion

Market Capitalization = Total Payout / [r(E) - g] = $5.33 billion / (0.128 - 0.084)

= $5.33 billion / 0.044 = $121.21 billion

b). P(0) = Market Capitalization / Number of Shares = $121.21 billion / 6.6 billion shares = $18.37